# On the Retention of Edited Knowledge in Fine-Tuned Language Models

### v1

@@ Abstract
Large language models (LLMs) store vast amounts of knowledge, which often requires updates to correct factual errors, incorporate newly acquired information, or adapt model behavior. Model editing methods have emerged as efficient solutions for such updates, offering localized and precise knowledge modification at significantly lower computational cost than continual training. In parallel, LLMs are frequently fine-tuned for a wide range of downstream tasks. However, the effect of fine-tuning on previously edited knowledge remains poorly understood. In this work, we systematically investigate how different fine-tuning objectives interact with various model editing techniques. Our findings show that edited knowledge is substantially more susceptible to forgetting during fine-tuning than intrinsic knowledge acquired through pre-training. This analysis highlights a key limitation of current editing approaches and suggests that evaluating edit robustness under downstream fine-tuning is critical for their practical deployment. We further find that freezing layers associated with edited content can significantly improve knowledge retention, offering insight into how future editing methods might be made more robust.
@@ 1 Introduction
LLMs have shown the ability to store vast amounts of knowledge  (radford2019a). The training of LLMs often involves with different stages, including pretraining and finetuning. Finetuning is important as it can align LLMs with human intent  (Ouyang2022follow). LLMs can acquire fundamental knowledge during the pretraining stage and subsequently adapt to specific tasks through downstream fine-tuning. However, when a gap exists between the pretraining and downstream finetuning stage, LLMs face the risk of catastrophic forgetting  (Wang2023).
As real-world information continually evolves, software must be constantly updated to keep pace to avoid becoming outdated  (Lazaridou2021). LLMs are simply another type of software that requires regular updates, particularly for knowledge refresh. Standard finetuning can bring large computation costs and perform poorly in locality and generalization  (Mitchell2022SERAC; Gangadhar2024ACL). To this end, the concept of knowledge editing (KE) has been proposed, enabling data-efficient modifications to the model behavior, while ensuring no adverse impact on other inputs. Recently, many methods have been proposed to correct the erroneous or obsolete knowledge  (Fang2024; Meng2023MEMIT; meng2023rome; Mitchell2022SERAC). The KE methods can be classified into 3 main categories: 1. Locate-then-edit 2. Meta-learning 3. Memory-based. Each category has some representing methods that proved to be efficient and effective.
Previous work has primarily evaluated KE methods based on reliability, portability, locality and efficiency in post-edit models  (Yao2023). However, the impact of downstream finetuning on edited knowledge remains poorly understood. In this paper, we address a critical question: how does downstream finetuning affect knowledge edited by KE methods, and how can we preserve them?
The downstream fine-tuning can also impair the model’s intrinsic knowledge  (Lange2022; Wu2022; Wang2023). To differentiate the finetuning impact on model edited knowledge and intrinsic knowledge, we create a dataset for each of them, and compare their retention rate of post-edit model and finetuned model. We conduct experiments on different KE methods-including standard finetuning, Locate-then-edit, and meta-learning method-and then finetune these post-edit models by downstream tasks including finetuning with unstructured and structured dataset, classification task as well as supervised finetuning(SFT) on multiple representative LLMs, including GPT-2 XL  radford2019a and Llama3-8B  (Dubey2024). We also make sure the finetuning corpus is irrelevant to the edit knowledge. We show different type of downstream finetuning tasks can affect differently to the intrinsic knowledge and edited knowledge. In addition, how the downstream finetuning task affect edited knowledge can depend on the edit method. An example is shown in Figure 1, even edit can be successful for some edit methods, the edited knowledge can be vulnerable to downstream fine-tuning tasks.
@@
Figure 1: Demonstration of model editing and downstream model finetuning and their impact on the knowledge in LLMs. Although the edit can be success, it is vulnerable to different downstream finetuning tasks. $f_{\theta}$, $f_{\theta^{\prime}}$, $f_{\theta"}$ denote the pretrained models, edited model and finetuned model respectively.
@@ 2 Related work
KE has emerged as a promising paradigm for updating LLMs to adapt to dynamically evolving information. There are plenty of works about KE methods, which can be classified into 3 main categories: 1. Locate-then-edit 2. Meta-learning 3. Memory-based.
Fine-tuning (FT). Vanilla FT is a straightforward method to modify the knowledge in LLMs  (Zhu2020FT). Model can be finetuned using the training corpus in the pretraining stage. Continual pre-training with a domain-specific corpus is an effective way to incorporate domain knowledge into a pre-trained model and enhance the its performance in that domain. However, the training corpus for fine-tuning often differs from the pre-training data. This can serve as a baseline for other model edit methods.
@@
Locate-then-edit. The locate-and-edit method for LLMs knowledge editing involves first identifying the specific parts of the model where the target knowledge is stored  (Dai2022; Meng2023MEMIT; meng2023rome). Once localized, the method directly modifies the model’s weights or representations in those areas to update or correct the knowledge. This approach aims to precisely edit knowledge while minimizing broader impacts on the model’s overall performance. ROME and MEMIT are examplars of this category. ROME edits factual knowledge in LLMs by identifying and updating specific rank-one subspaces in the model’s weights, allowing precise, localized changes without retraining  (meng2023rome). It leverages causal tracing to locate key layers and modifies them efficiently to correct or update facts while preserving the model’s overall performance.
@@
Meta-learning Method. Meta-learning for LLMs knowledge editing involves training a hyper-network to generate targeted parameter shifts that update the model’s knowledge without full retraining  Mitchell2022MEND; Cao2022knowledge. This approach leverages the hyper-network to transform standard fine-tuning gradients into precise edits, ensuring generalization to semantically equivalent inputs while preserving unrelated knowledge. MALMEN is a representative of the meta-learning-based editing approach. MALMEN edits large language models by using a hyper-network to compute parameter shifts as a least squares problem, solved via the normal equation, enabling efficient and scalable updates while minimizing interference with unrelated knowledge.
Memory-based Methods. Memory-based methods store edits in a explicit memory without modifying the model’s parameters  Mitchell2022SERAC; Zheng2023ACL; Hartvigsen2023. For example, SERAC is a gradient-free memory-based model editing method that stores edits in an explicit memory and uses a scope classifier to determine if a test input is within the scope of any cached edits. If within scope, a counterfactual model predicts the label based on the most relevant edit example; otherwise, the base model’s prediction is used  Mitchell2022SERAC.
@@
Our work focuses specifically on these the first two parameter-modifying approaches, as they enable more fundamental alterations to the model’s knowledge representations. Some work has show that KE methods fail to retain the model’s accuracy on irrelevant knowledge and general ability  Gupta2024; Gu2024, and the evaluation paradigm for model editing has been investigated  (Cohen2024; Zhong2023). Besides, recent work reveals significant limitations in current methods, showing that their performance on real-world hallucinations often falls short of expectations, and highlights the need for further improvements in the field  (huang2024).
3 Retention analysis of edited knowledge after fine-tuning
@@ 3.1 Experiment setup
We evaluate whether a language model encodes specific knowledge using COUNTERFACT, a dataset contains counterfactual statements, is designed to distinguish superficial lexical changes from meaningful alterations in factual knowledge. In COUNTERFACT, each record is knowledge in a simple form that contains a subject, a relation and a object  (meng2023rome). Each COUNTERFACT instance is derived from PARAREL and consists of a true knowledge tuple $t^{c}=(s,r,o^{c})$ and a false knowledge tuple $t^{*}=(s,r,o^{*})$, where s stands for subject, r stands for relation, $o^{c}$ stands for correct object and $o^{*}$ stands for false object. The subject and relation can form a prompt. Edit methods in this paper are performed to edit the corresponding object. We create our edited knowledge dataset from COUNTERFACT train split and intrinsic knowledge dataset from the test split.
Edited knowledge dataset The edited knowledge dataset is constructed by sampling from the model’s existing knowledge. To achieve this, we filter the dataset by retaining only instances that given the prompt $p=(s,r)$, the model assigns the true target token with the highest probability that significantly surpasses that of other tokens by a substantial margin. From this filtered set, we sample 50 instances to form the edited knowledge dataset.
Intrinsic knowledge dataset We create the intrinsic knowledge dataset for benchmark purpose. Similar to the edited knowledge dataset, the intrinsic knowledge dataset is crafted by sampling 100 data that the true target token is assigned with the significant highest probability. We make sure that there is no overlap between intrinsic knowledge dataset and edited knowledge dataset for the subject, relation as well as object. More details about the dataset are shown in Appendix A.
Knowledge editing This study investigates three distinct KE methods: FT, ROME and MALMEN, that representing three categories of approaches: 1) Brutal-force Fine-tuning, 2) Locate-then-Edit, and 3) Meta-learning respectively. To enhance locality, we fine-tune only a specific layer while freezing all others, as this approach demonstrates superior locality compared to full-model fine-tuning  (Gangadhar2024ACL). A study has been conducted to determine the optimal layer for knowledge edits  (Hase2024NIPS; meng2023rome). For FT and ROME method, we choose layer 1 and layer 6 to be the edit layer respectively. In all of these experiments, including FT, ROME and MALMEN, we stop editing until the loss goes below a threshold. As a result, we achieve perfect edit success rate.
Downstream fine-tuning We conduct four different type of downstream fine-tuning tasks 1) Fine-tuning with unstructured text, 2) Fine-tuning with structured factual text, 3) Classification tasks, and 4) Question-answer supervised finetuning task on GPT-2 XL and Question-answer supervised finetuning on Llama3-8B. To fairly compare these edit methods, it is essential to develop a metric to quantify the extent to which fine-tuning influences the model. To ensure that the fine-tuning impact is consistent across the three methods (FT, ROME, and MALMEN), for each of these downstream fine-tuning tasks, we sample and fix an evaluation dataset from the fine-tuning dataset and standardize the training stopping criteria, and employ the same finetuning hyperparameter.
  • Fine-tuning with the Unstructured Dataset To systematically assess the impact on both the model’s intrinsic knowledge and the edited knowledge, we employed an unstructured dataset distinct from all pre-training corpora. Specifically, for our experiments on the GPT-2 XL, which was pre-trained on webtext, we choose the Common Crawl dataset for fine-tuning. Since a small subset of the data suffices to demonstrate the influence, we sampled 60k data for our training set. We evaluate the finetuning influence on a validation set, which contains of 1k instances sampled from the training set. We employ loss as a metric to measure the volume of knowledge acquired from fine-tuning. The finetuning is stopped once the validation loss goes below the threshold of 3.1.
  • Finetuning with the Structured Factual Dataset To investigate the influence of structured factual data on model behavior, we construct a fine-tuning dataset that mirrors the structure of our edit data while containing distinct factual content. Specifically, we sample data from COUNTERFACT train split, which has no overlap with the one used in upstream edit. We select only instances where the pretrained model fails to predict the true target, and sample 3,000 of them. For supervision, we use the true target from each example as the training label. To monitor progress, we evaluate model performance on a validation set, which contains 100 instances sampled from the training set. Training terminates once the model achieves above 70% accuracy on this validation set.
  • Classification Task To assess the impact of downstream fine-tuning on pretrained knowledge, we employ a classification task as our benchmark evaluation. Specifically, we utilize the IMDB sentiment analysis dataset which consists of 25k movie reviews paired with binary sentiment labels. For the classification architecture, we append a fully connected layer to the final hidden state of the end-of-sequence (EOS) token. Classification accuracy on the validation set can be a fair metric, as it is the target index people are aiming to improve. The validation set contains 100 samples, which has no overlap with the training set. We finetune the model until the model achieves a classification accuracy exceeding 95% on the validation set.
  • Supervised Finetuning Task Supervised fine-tuning has emerged as a prevalent downstream adaptation method for LLMs. In this paradigm, each training instance consists of an instruction (question) paired with its corresponding response (answer). For our experiments, we employ the Tulu-3-SFT-mixture dataset  Lambert2024tulu for instruction fine-tuning, which consists of 10k movie reviews paired with binary sentiment labels. To assess the training influence, we evaluate model performance on a validation set, which contains 1k instances sampled from the training set. Training terminates once the model achieves above 70% accuracy on this validation set. We evaluate the training progress by compute the model’s loss on the validation set. We set the stopping criteria as validation loss goes below the threshold of 2.5.
@@
To assess whether the model possesses the target knowledge, we input the prompt into the model and compute the probability distribution over the output token. The model can sometimes assigns the highest probability to a stop word but not the target word. However, this does not necessarily imply a lack of knowledge. In such cases, we employ greedy decoding to generate subsequent tokens until a non-stop token is produced. We consider the model to retain the knowledge if the first non-stop token matches the target token. We define the edited knowledge retention rate as the proportion of prompts for which the target token is the highest-ranked non-stop word:
@@
Where $y_{t}$ is the target token, $\{(x,y)\}$ is our edited/intrinsic dataset, S is the set of stopping word first token. We can obtain a intrinsic knowledge retention rate similar to this. We evaluating both the edited knowledge retention rate and intrinsic knowledge retention rate on both post-edit and finetuned model. Details about these model edit and finetuning hyper-parameters are shown in Appendix B.
For a combination of edit method and downstream finetuning task, we:
  • For each knowledge tuple from the edited knowledge dataset, we conduct a single-edit on the original model to obtain a post-edit model using this edit method
  • Fine-tuning the post-edit model on the downstream tasks
  • Assess the edited knowledge and intrinsic knowledge retention rate for both post-edit model and finetuned knowledge on the single edited knowledge and intrinsic knowledge dataset.
3.2 Finetuning impact on edited and intrinsic knowledge
Figure 2: Edited and intrinsic knowledge retention rate after model edit and finetuning for different combination of upstream edit methods and downstream finetuning methods. For ROME method, we choose layer 6 as the edit layer. For FT method, we choose layer 1 as the edit layer.

### v2

@@ Abstract
Large language models (LLMs) store vast amounts of knowledge, which often requires updates to correct factual errors, incorporate newly acquired information, or adapt model behavior. Model editing methods have emerged as efficient solutions for such updates, offering localized and precise knowledge modification at significantly lower computational cost than continual training. In parallel, LLMs are frequently fine-tuned for a wide range of downstream tasks. However, the effect of fine-tuning on previously edited knowledge remains poorly understood. In this work, we systematically investigate how different fine-tuning objectives interact with various model editing techniques. Our findings show that edited knowledge is substantially more susceptible to forgetting during fine-tuning than intrinsic knowledge acquired through pre-training. This analysis highlights a key limitation of current editing approaches and suggests that evaluating edit robustness under downstream fine-tuning is critical for their practical deployment. We further find that knowledge retention can be significantly improved by either augmenting edit knowledge with paraphrases or by freezing layers associated with edited content in fine-tuning stage, offering insight for developing more robust editing algorithms.
On the Retention of Edited Knowledge in Fine-tuned Language Models
Fufang Wen Columbia University fw2325@columbia.edu          Shichang Zhangthanks: Corresponding author Harvard University shzhang@hbs.edu
@@ 1 Introduction
A key factor in the success of large language models (LLMs) is their ability to store vast amounts of knowledge (Radford et al., 2019). The stored knowledge serves as a foundation that enables LLMs to be readily adapted to specific downstream tasks or aligned with human intent through fine-tuning (Ouyang et al., 2022). Fine-tuning has become an essential step in LLM development, with the majority of production models undergoing this process as developers adapt base models for various applications. However, the mechanisms regarding knowledge updates during fine-tuning remain unclear, and the risk of catastrophic forgetting persists (Lange et al., 2022; Wu et al., 2022; Wang et al., 2023). As fine-tuning becomes increasingly sophisticated and necessary to LLM development, understanding and improving fine-tuning knowledge retention has emerged as a critical challenge.
While LLMs acquire most of their knowledge through pre-training on large corpora, they can also be updated through direct knowledge editing (KE). The KE procedure is analogous to software maintenance: as traditional software requires bug fixes to maintain functionality, LLMs need mechanisms to update their knowledge when it is incorrect or becomes outdated. For example, when a country elects a new president, the LLM’s knowledge regarding the country-president relationship needs to be updated accordingly. Though such updates could be implemented through continued training, this approach is both costly and risks overfitting Mitchell et al. (2022b); Gangadhar and Stratos (2024). KE methods provide more data-efficient and precise solutions for implementing these updates without full model retraining (Fang et al., 2024; Meng et al., 2023b, a; Mitchell et al., 2022b). However, a crucial question remains: does edited knowledge retain during subsequent fine-tuning stages as effectively as intrinsic knowledge learned from pre-training? The retention of edited knowledge through fine-tuning is crucial for LLM development, analogous to ensuring that fixed software bugs do not reappear in later development cycles.
In this paper, we start by conducting systematic studies on applying different KE methods on LLMs and evaluating the knowledge retention after fine-tuning, as illustrated in Figure 1. We consider four different types of fine-tuning tasks: next token prediction on general text, object prediction for factual knowledge triplets, sentiment classification, and instruction fine-tuning. We find that while the knowledge retention rate depends on both the editing method and the task, it is in general highly vulnerable to fine-tuning, and the retention rate of the edited knowledge is significantly lower than that of intrinsic knowledge.
We then connect the knowledge retention with the elasticity theory of LLM fine-tuning (Ji et al., 2025) to explain this fragility. This theory suggests that the LLM’s resistance to deviate from its original distribution is proportional to the training data volume. According to this theory, since the intrinsic knowledge is trained on vastly more data, it achieves greater resistance against fine-tuning than edited knowledge. Therefore, to bridge the gap of retention edited knowledge, we propose two strategies to enhance the resilience of edited knowledge. One is editing via augmenting paraphrases. Experimental results confirm that this data augmentation significantly improves retention, matching or even surpassing the retention rate of intrinsic knowledge when sufficient paraphrases are provided. The other is fine-tuning via freezing layers. We show that selectively freezing layers most associated with the edit content during downstream training effectively preserves edits, achieving retention rates on par with intrinsic knowledge and confirming the localizability of edited knowledge.
The primary contributions of this work are summarized as follows:
  • To the best of our knowledge, we conduct the first systematic study on the retention of edited knowledge in LLMs after fine-tuning. We demonstrate that successfully edited knowledge remains highly vulnerable to fine-tuning across a range of editing methods and fine-tuning tasks.
  • We connect the knowledge retention with model elasticity theory to explain the low retention rate of the edited knowledge. In the light shed by this theory, we propose and validate two effective strategies for improving edit resilience. By paraphrase-augmented editing, we expand edit instances with multiple paraphrases of the target fact. By fine-tuning via freezing layers, we selectively freezing layers most associated with the edit content during downstream training.
@@
Figure 1: Demonstration of model editing and downstream model fine-tuning and their impact on the knowledge in LLMs. The original model is edited with a single instance of new fact: Windows Mobile 6.5 was developed by Apple, and the edited model is fine-tuned by an irrelevant dataset, which does not contain subject, relation and object from the edited knowledge. Although the edit can be successful, it is vulnerable to different downstream fine-tuning tasks. $f_{\theta}$, $f_{\theta^{\prime}}$, $f_{\theta"}$ denote the pre-trained models, edited model and fine-tuned model respectively.
@@ 2 Related work
KE has emerged as a promising paradigm for updating LLMs to adapt to dynamically evolving information. Vanilla fine-tuning is a straightforward method to modify the knowledge in LLMs  Zhu et al. (2020). There are plenty of works about KE methods, which can be classified into 3 main categories: 1. Locate-then-edit 2. Meta-learning 3. Memory-based.
@@
Locate-then-edit. The locate-and-edit method for LLMs knowledge editing involves first identifying the specific parts of the model where the target knowledge is stored  (Dai et al., 2022; Meng et al., 2023b, a). Once localized, the method directly modifies the model’s weights or representations in those areas to update or correct the knowledge. This approach aims to precisely edit knowledge while minimizing broader impacts on the model’s overall performance. ROME and MEMIT are examplars of this category. ROME edits factual knowledge in LLMs by identifying and updating specific rank-one subspaces in the model’s weights, allowing precise, localized changes without retraining  (Meng et al., 2023a). It leverages causal tracing to locate key layers and modifies them efficiently to correct or update facts while preserving the model’s overall performance. Building on ROME, MEMIT is a method for mass-editing factual associations in LLMs by directly updating the weights of specific transformer MLP layers identified as causal mediators for factual recall. It spreads the desired memory updates across multiple layers.
@@
Meta-learning for KE. Meta-learning for LLMs knowledge editing involves training a hyper-network to generate targeted parameter shifts that update the model’s knowledge without full retraining  (Mitchell et al., 2022; Cao et al., 2022). This approach leverages the hyper-network to transform standard fine-tuning gradients into precise edits, ensuring generalization to semantically equivalent inputs while preserving unrelated knowledge. MALMEN is a representative of the meta-learning-based editing approach. MALMEN edits large language models by using a hyper-network to compute parameter shifts as a least squares problem, solved via the normal equation, enabling efficient and scalable updates while minimizing interference with unrelated knowledge.
@@
Some work has show that KE methods fail to retain the model’s accuracy on irrelevant knowledge and general ability Gupta et al. (2024); Gu et al. (2024), and the evaluation paradigm for model editing has been investigated  (Cohen et al., 2024; Zhong et al., 2023). Besides, recent work reveals significant limitations in current methods, showing that their performance on real-world hallucinations often falls short of expectations, and highlights the need for further improvements in the field  (Huang et al., 2025).
Model Elasticity Theory. Recent work has begun to theorize the underlying mechanisms of alignment fragility. (Ji et al., 2025) propose the "elasticity" theory, arguing from a data compression  (Delétang et al., 2023) perspective that language models inherently resist distributional shifts from alignment fine-tuning due to the overwhelming influence of pre-training data, causing them to easily revert to pre-trained behaviors.
3 Knowledge Retention Analysis for Fine-Tuned LLMs
We investigate the retention of both edited and intrinsic knowledge in LLMs after fine-tuning. We evaluate a range of KE methods and fine-tuning tasks under the pipeline: beginning with a base model possessing intrinsic knowledge, we apply a KE technique to introduce edited facts, perform fine-tuning, and subsequently assess the retention of both knowledge types. Here, "knowledge" refers to the model’s ability to correctly answer queries regarding a given subject-relation-object triple. They are in the form of $t^{c}=(s,r,o^{c}/o^{*})$, where s stands for subject, r stands for relation, $o^{c}$ stands for correct object and $o^{*}$ stands for false object. The subject and relation together form a prompt used to query the model. It can be evaluated by giving the model a prompt of (subject, relation) and checking if the model can generate the correct object.
Each instance in COUNTERFACT, which is derived from PARAREL, In this work, KE methods are applied to modify the model’s association for a given $(s,r)$ pair, replacing the original object $o^{c}$ with the new, edited object $o^{*}$ pair.
@@ 3.1 Experiment Setup
KE Dataset: We constructed a dataset to evaluate the retention of both edited and intrinsic knowledge after fine-tuning. Both types of knowledge were derived from the model’s existing knowledge base. We use the COUNTERFACT dataset Lambert et al. (2024) as the base dataset because it contains factual knowledge in the form of (subject, relation, object) triples as well as the corresponding counterfactual object. We treat the counterfactual object as the edited knowledge, and the factual object as the intrinsic knowledge. For clean evaluation, we further filter the COUNTERFACT dataset and retain only instances where, given a prompt $p=(s,r)$, the model assigned the highest probability to the true target token by a substantial margin. Furthermore, we ensured that there was no overlap between the intrinsic knowledge dataset and the edit knowledge dataset in terms of subjects, relations, and objects. Further details regarding the dataset are provided in Appendix B.
KE Methods: We investigate five distinct KE methods: A baseline method performing full fine-tuning for editing(FT), ROME, MEMIT, MALMEN and AlphaEdit Meng et al. (2023a, b); Tan et al. (2023); Fang et al. (2024). For single-edit method (e.g., ROME), the model is updated with one knowledge tuple at a time. For the batch-edit method (e.g., MEMIT), the knowledge dataset is partitioned into groups containing multiple tuples, and the model is updated with an entire group simultaneously to produce the post-edit model. For FT and ROME, we also try different edited layer. We find that for both of the methods, edit early layer (prior to layer 10) can achieve similar result. More details about KE methods are shown in Appendix C.1.
Fine-Tuning Tasks and Datasets: We fine-tune the post-edit model on four different downstream tasks: 1) Next token prediction on general text in CommonCrawl, 2) Object prediction for factual knowledge triplets in COUNTERFACT, 3) Sentiment classification in IMDB, and 4) Instruction fine-tuning in Tulu-3 SFT Lambert et al. (2024). To prevent potential conflicts between the fine-tuning data and edited/intrinsic knowledge—where the former may contain contradictory information that could degrade the edited knowledge—we systematically filter the fine-tuning dataset. To ensure fairness of experiment, we set the same stopping criteria for fine-tuning different post-edit models. More details about the fine-tuning tasks are shown in Appendix C.2.
Figure 2: Edited and intrinsic knowledge retention rate after model edit and fine-tuning for different combination of upstream edit methods and downstream fine-tuning methods. For ROME method, we choose layer 6 as the edit layer. For FT method, we choose layer 1 as the edit layer.
Models: We perform the KE and fine-tuning on the GPT-2 XL Radford et al. (2019) and Llama3-8B (Dubey et al., 2024).
@@
Evaluation: We evaluate the edit and intrinsic knowledge retention rate for both the post-edit model and the fine-tuned model. To assess whether the model possesses the target knowledge, we input the prompt into the model and compute the probability distribution over the output token. We consider the model to retain the knowledge if the first non-stop token matches the target token. We define the knowledge retention rate as the proportion of prompts for which the target token is the highest-ranked non-stop word:
@@
Where $y_{t}$ is the target token, $\{(x,y)\}$ is our edited/intrinsic dataset, S is the set of stopping word first token. We can obtain a intrinsic knowledge retention rate similar to this. We evaluate both the edited knowledge retention rate and intrinsic knowledge retention rate on both post-edit and fine-tuned model. Details about these model edit and fine-tuning hyper-parameters are shown in Appendix C.2.
3.2 Knowledge Retention After Fine-tuning